# Quantum framework for Reinforcement Learning: integrating Markov Decision Process, quantum arithmetic, and trajectory search

### v1

@@ Abstract
This paper introduces a quantum framework for addressing reinforcement learning (RL) tasks, grounded in the quantum principles and leveraging a fully quantum model of the classical Markov Decision Process (MDP). By employing quantum concepts and a quantum search algorithm, this work presents the implementation and optimization of the agent-environment interactions entirely within the quantum domain, eliminating reliance on classical computations. Key contributions include the quantum-based state transitions, return calculation, and trajectory search mechanism that utilize quantum principles to demonstrate the realization of RL processes through quantum phenomena. The implementation emphasizes the fundamental role of quantum superposition in enhancing computational efficiency for RL tasks. Experimental results demonstrate the capacity of a quantum model to achieve quantum advantage in RL, highlighting the potential of fully quantum implementations in decision-making tasks. This work not only underscores the applicability of quantum computing in machine learning but also contributes the field of quantum reinforcement learning (QRL) by offering a robust framework for understanding and exploiting quantum computing in RL systems.
@@ I Introduction
This paper is organized as follows: Sec. II reviews previous studies on integrating RL with QC. Section III introduces key concepts in classical reinforcement learning and quantum computing. Section IV details our quantum implementation, including circuit designs, quantum interactions, and trajectory search. Section V presents experimental results, and Sec. VI and Sec. VII summarize insights and propose future research directions.
@@ II.4 Full quantum reinforcement learning
In our work, we aim to address the limitations that previous works have encountered. Methods employing VQCs encoded only the agent as a quantum component while retaining the environment in the classical domain, restricting the quantum interactions of agent and environment. Other approaches, such as QiRL and quantum subroutines, relied on classical computations, preventing the realization of fully quantum operations. Furthermore, prior work on a fully quantum system [35] was constrained to a single-state bandit setting, which does not generalize to complex decision-making scenarios involving finite-state Markov Decision Processes (MDPs).
In contrast, our work extends to a finite-state MDP by introducing a fully QRL framework, where both the agent and the environment are encoded as quantum components and their interactions occur entirely within the quantum domain. The concept of a quantum-accessible environment was first introduced in [36] where the agent and environment interact through an intermediate register. In comparison, our method enables direct communication between the agent and environment, eliminating the need for an additional layer. This direct interaction allows for the simultaneous evaluation of multiple states and actions, increasing computational parallelism and enabling faster exploration of the state-action space. Furthermore, since these interactions are entirely within the quantum domain, this will remove the classical-quantum conversions and achieve the complete quantum interactions. Our framework facilitates optimization by a quantum search algorithm directly to the quantum samples generated through these interactions. This approach eliminates reliance on classical subroutines and achieve significant computational speedups. To the best of our knowledge, our work is the first attempt to implement a complete quantum framework for reinforcement learning by integrating the quantum realization of the classical Markov Decision Process (MDP).
@@ III.2 Quantum computing
Quantum computing (QC) combines quantum mechanics with computer science to enable advanced computation and information processing [39]. In QC, information is represented using qubits, which can exist in superposition states. This unique property allows qubits to process vast amounts of information simultaneously. The quantum computation process consists of three primary steps. The first step is state initialization, where qubits are typically prepared into uniform superposition. The second step involves the application of quantum gates to perform the computation. These gates are represented by unitary matrices, ensuring that transformations applied to quantum states are reversible and that the total probability is preserved [40]. The final step is measurement, where qubits collapse from their quantum superposition into classical states, producing probabilistic outcomes that represent the results of the computation. These steps collectively enable the construction of complex quantum algorithms capable of solving advanced problems such as Shor’s algorithm for integer factorization [41, 42, 43] and Grover’s algorithm for searching unsorted datasets [12]. An example of a quantum circuit for Grover’s algorithm is shown in Fig. 2, illustrating the process from initialization and computation (oracle and diffuser operations) to the final measurement stage.
Figure 2: Quantum circuit for Grover’s algorithm on 2 qubits, searching for the target state |$\ket{11}$. The measurement was performed using IBM quantum processor $(ibm\_brisbane)$. Output distribution is displayed on the right, showing the search state |$\ket{11}$ with the highest count.
@@ IV.1 Quantum implementation of Markov Decision Process
In reinforcement learning, the agent interacts with the environment within a Markov Decision Process (MDP) framework. The agent selects actions based on a policy, which maps states to actions, while the environment determines state transitions and provides rewards. This interaction is cyclic: the agent observes a state, takes an action, receives a reward, and updates its policy to maximize cumulative rewards over time. Key MDP components include states, actions, a state transition function, and a reward function. For the MDP framework, we consider a simple stochastic environment with four states $(s0,s1,s2,s3)$ and two actions $(a0,a1)$, where the agent moves from one state to another with certain probabilities depending on the chosen action. For example, if the agent is in state $s0$ and selects action $a1$, there is some probability that it will move to $s1$, and some probability it will remain in $s0$ or will move to another state. This probability of transition from one state to another is determined by the state transition function. The reward function assigns a numerical value to each transition, representing the benefit of moving from one state to another. For instance, transitioning to state $s1$ might yield a reward $r1$. The goal of the agent is to find the optimal policy, a strategy that determines which action to take in each state. The optimal policy aims to maximize the total reward the agent can collect over time. Classical Q-learning solves this problem by allowing the agent to learn from its interactions with the environment. The agent explores the states, takes actions, and observes the resulting rewards and transitions. It builds a table, known as a Q-table, that estimates the expected cumulative reward for each state-action pair. By balancing exploration (trying new actions) and exploitation (choosing actions with the highest known rewards), the agent identifies the optimal actions for each state.
@@ IV.1.1 Agent and environment on quantum computer
As we are implementing MDP in the quantum domain, these states and actions can be prepared into superposition by using Hadamard transformation. This process can initialize qubits as a uniform combination of quantum samples for all states and actions within their respective spaces. For the state space S, a Hadamard operator transforms the initial ground state qubit |$\ket{0_{s}}$ into the uniform superposition quantum states, formulated as in Eq. (2).
$H\ket{0_{s}}=\sum_{n=1}^{N}c_{n}\ket{s_{n}}.$ (2)
The output quantum state is a uniform superposition of the basis states, |$\ket{s_{1}},\ket{s_{2}},...,\ket{s_{n}}$, within its state space S. N represents the total number of states in that space and the coefficient $c_{n}$ is the amplitude of each quantum state, which matches the normalization condition of the valid quantum state, $\sum_{n=1}^{N}|c_{n}|^{2}=1$.
Similarly, the set of possible actions is encoded using the Hadamard operator to create a uniform superposition of all actions which belong to the action space A, as formalized in Eq. (3).
$H\ket{0_{a}}=\sum_{n=1}^{N}c_{n}\ket{a_{n}}.$ (3)
@@
After preparing the state S and action A into uniform superposition quantum samples, we then consider the state transition function, which defines the likelihood of transitioning from one state to another given a specific action. In a classical MDP, the state transition function $P(s^{\prime}|s,a)$ determines the probability of transitioning to next state ${s^{\prime}}$ given the current state $s$ and action $a$. In a quantum circuit, these probabilities are encoded in the amplitudes of quantum states, where the squared magnitudes of the amplitudes correspond to probabilities. This is achieved using $R_{y}(\theta)$ gates, which rotate a qubit around the Y-axis of the Bloch sphere, producing a superposition where $\cos^{2}\left(\frac{\theta}{2}\right)\text{ and }\sin^{2}\left(\frac{\theta}{2}\right)$ represent the probabilities of measuring |$\ket{0}$ and |$\ket{1}$, respectively. To encode classical transition probabilities $P(s^{\prime}|s,a)$, the rotation angle $\theta$ is calculated as $\theta=2\arcsin(\sqrt{P(s^{\prime}|s,a)})$. In our work, we use multi-controlled $R_{y}(\theta)$ gates to apply these rotations conditionally based on the current state |$\ket{s}$ and action |$\ket{a}$. The rotation is applied only when the control qubits match a specific state-action pair. Equation (6) is a mathematically representation of this quantum state transition process where the controlled $R_{y}(\theta)$ gates apply a rotation to an ancillary ground state qubit |$\ket{0_{s^{\prime}}}$ by an angle $\theta$, depending on the state-action pair, thereby determining the next state, |$\ket{s^{\prime}}\in S$.
@@ IV.2 Agent-environment quantum interactions over multiple time steps
To extend the QMDP framework over H time steps, we need to account for the quantum state transition and the calculation of cumulative rewards. At each time step t, the agent receives a quantum state |$\ket{s_{t}}$, performs an action |$\ket{a_{t}}$, and the environment responds by transitioning to a new state |$\ket{s^{\prime}_{t}}$ while providing a reward |$\ket{r_{t}}$. This process is repeated over H time steps, ensuring that the next state |$\ket{s^{\prime}_{t}}$ at t becomes the current state |$\ket{s_{t}}$ for the subsequent time step.
@@
Below is a description of how this process unfolds over $H$ time steps, including the corresponding quantum operators that govern state transitions and rewards. The approach builds on the single interaction discussed earlier and extends it to a sequence of interactions. Initially at time step ${t=0}$, the agent starts at state |$\ket{s_{0}}$, takes an action |$\ket{a_{0}}$, transition to the next state |$\ket{s^{\prime}_{0}}$ and receive a reward |$\ket{r_{0}}$. The next state |$\ket{s^{\prime}_{0}}$ propagates to the subsequent time step. At time step ${t=1}$, the agent observes the new state |$\ket{s_{1}}$ generated at ${t=0}$, selects a new action |$\ket{a_{1}}$, and the environment transitions to the next state |$\ket{s^{\prime}_{1}}$ and provides the reward |$\ket{r_{1}}$. Once again, the next state |$\ket{s^{\prime}_{1}}$ becomes the current state for the subsequent time step ${t=2}$. In this way, the interaction process will occur for $H$ time steps. The quantum state distribution at ${t=1}$ can be expressed as in Eq. (9), where $N$ represents the total number of trajectories in that quantum interaction. The terms $c_{s^{\prime}_{0},r_{0}|s_{0},a_{0}}$ and $c_{s^{\prime}_{1},r_{1}|s_{1},a_{1}}$ are the amplitudes indicating that the system transitions to the next state |$\ket{s^{\prime}}$ with reward |$\ket{r}$, given current state |$\ket{s}$ and action |$\ket{a}$ at time step ${t=0}$ and ${t=1}$, respectively.
@@
For the $H$ time steps of interaction, each quantum sample can be represented as a string of qubits encoding the possible transitions from ${t=0}$ to ${t=H-1}$, formed by concatenating state-action-next state-reward sequences. This is obtained by applying a sequence of individual interactions. The complete quantum distribution of the agent-environment interaction at time step $H$ can be summarized as Eq. (10).
|$\ket{\phi^{(H)}}=\sum_{n=1}^{N}\bigg{[}\sum_{t=0}^{H-1}c_{s^{\prime}_{t},r_{t}|s_{t},a_{t}}\prod_{t=0}^{H-1}\ket{s_{t}}\ket{a_{t}}\ket{s^{\prime}_{t}}\ket{r_{t}}\bigg{]}^{n}.$ (10)
The entire quantum state after $H$ time steps is a quantum superposition over all possible trajectories $N$, where each trajectory consists of state, action, next state and reward sequences from ${t=0}$ to ${t=H-1}$. The amplitudes $\sum_{t=0}^{H-1}c_{s^{\prime}_{t},r_{t}|s_{t},a_{t}}$ define the likelihood of each trajectory based on the agent’s actions and environment’s responses up to time step $H$.
@@ IV.2.1 Return calculation by quantum arithmetic
In our work, this classical concept is mapped into the quantum domain, where the return is encoded as a quantum state |$\ket{g}$ in a Hilbert space $G$. This space represents a quantum system with a sufficient number of qubits to encode every potential return value. Initially, the quantum state |$\ket{g}$ is prepared in its ground state |$\ket{0_{g}}$. To compute return, unitary operation $U_{G}$ is applied to the reward qubits |$\ket{r_{t}}$ for each time step. The unitary operation $U_{G}$ executes a quantum addition that performs sequential bit-wise addition of each reward register |$\ket{r_{t}}$ and stores the result in the return registers |$\ket{g}$. This operation propagates carry bits through the circuit, similar to a classical addition. For each reward register |$\ket{r_{t}}$, a series of controlled-NOT and Toffoli gates are applied to implement bit-wise addition into the return registers. This unitary operation $U_{G}$ can be described as Eq. (11), where $\gamma\in[0,1]$ is a fixed discount factor, $H$ is the total number of time steps, and |$\ket{r_{t}}$ represents the reward at time step $t$.
$U_{G}:=\sum_{t=0}^{H-1}\gamma\ket{r_{t}}.$ (11)
@@
When applying the return operation $U_{G}$ to the Quantum Markov Decision Process (QMDP) over $H$ time steps, the goal is to compute the return for each possible trajectory by summing the discounted rewards across those time steps. The final quantum state over these interactions can be expressed as Eq. (12).
@@
where N is the total number of possible quantum trajectories for the agent-environment interactions over $H$ time steps, |$\ket{g}$ represents the return for each trajectory, which sums the discounted rewards over the $H$ time steps. Therefore, the complete quantum state |$\ket{\phi^{(H)}}$ contains not only the state, action, next state and reward sequences but also the return for each trajectory. This enables the QMDP to evaluate the performance of all possible sequences of interactions efficiently by computing the return across multiple trajectories simultaneously.
The quantum state |$\ket{\phi^{(H)}}$, discussed in Eq. (12), is not human-interpretable classical information. The coefficients $c_{s^{\prime}_{t},r_{t}|s_{t},a_{t}}$ that define the amplitude of the quantum state cannot be accessed directly due to the nature of quantum systems. The only way to obtain the information from the quantum state |$\ket{\phi^{(H)}}$ is by measurement. Measurement is a probabilistic process that results a specific quantum state with a likelihood determined by $|c_{s^{\prime}_{t},r_{t}|s_{t},a_{t}}|^{2}$ Therefore, the quantum state |$\ket{\phi^{(H)}}$ can be regarded as quantum encoding of probability distributions over the possible trajectories of the agent-environment interactions across $H$ time steps.

### v2

@@ Abstract
This paper introduces a quantum framework for addressing reinforcement learning (RL) tasks, grounded in the quantum principles and leveraging a fully quantum model of the classical Markov Decision Process (MDP). By employing quantum concepts and a quantum search algorithm, this work presents the implementation and optimization of the agent-environment interactions entirely within the quantum domain, eliminating reliance on classical computations. Key contributions include the quantum-based state transitions, return calculation, and trajectory search mechanism that utilize quantum principles to demonstrate the realization of RL processes through quantum phenomena. The implementation emphasizes the fundamental role of quantum superposition in enhancing computational efficiency for RL tasks. Results demonstrate the capacity of a quantum model to achieve quantum enhancement in RL, highlighting the potential of fully quantum implementations in decision-making tasks. This work not only underscores the applicability of quantum computing in machine learning but also contributes the field of quantum reinforcement learning (QRL) by offering a robust framework for understanding and exploiting quantum computing in RL systems.
@@ I Introduction
This paper is organized as follows: Sec. II reviews previous studies on integrating RL with QC. Section III introduces key concepts in classical reinforcement learning and quantum computing. Section IV details our quantum implementation, including circuit designs, quantum interactions, and trajectory search. Section V presents demonstrations and their results, and Sec. VI and Sec. VII summarize insights and propose future research directions.
@@ II.4 Full quantum reinforcement learning
In our work, we aim to address the limitations that previous works have encountered. Methods employing VQCs encoded only the agent as a quantum component while retaining the environment in the classical domain, restricting the quantum interactions of agent and environment. Other approaches, such as QiRL and quantum subroutines, relied on classical computations, preventing the realization of fully quantum operations. Furthermore, while prior work on a fully quantum system [35] focused on a single-state bandit setting with only one state, our framework extends to more complex decision-making scenarios by implementing a finite-state Markov Decision Process (MDP) with multiple states, actions, and stochastic transitions. This generalization makes our model applicable to a broader range of practical reinforcement learning problems beyond simple bandit-like tasks. Our framework encodes both the agent and the environment as quantum components, enabling sequential agent-environment interactions across multiple time steps. It incorporates quantum state transitions and return calculations using quantum arithmetic. Our approach enables the quantum agent to explore numerous sequences of interactions simultaneously. Such parallelism enables efficient exploration of trajectories, allowing the agent to evaluate multiple potential outcomes concurrently and accelerate the discovery of high-reward policies. In contrast to this prior work, which employed Grover’s search to identify high-value actions in a static, single-step scenario, our framework applies Grover’s algorithm to search for an optimal trajectory among all possible full-length quantum trajectories in a multi-state, multi-step MDP. By designing a quantum oracle that marks trajectories based on their cumulative return, our approach enables direct optimization of agent’s performance over multiple time steps. This significantly expands the applicability of Grover’s search beyond action-level selection, enabling efficient discovery of optimal policies in multi-step reinforcement learning tasks.
To further advance fully quantum RL, our framework enables direct interaction between the quantum agent and the environment entirely within the quantum domain. The concept of a quantum-accessible environment was first introduced in [36] where the agent and environment interact through an intermediate register. In comparison, our method enables direct communication between the agent and environment, eliminating the need for an additional layer. This direct interaction allows for the simultaneous evaluation of multiple states and actions, increasing computational parallelism and enabling faster exploration of the state-action space. Furthermore, since these interactions are entirely within the quantum domain, this will remove the classical-quantum conversions and achieve the complete quantum interactions. Our framework facilitates optimization by a quantum search algorithm directly to the quantum samples generated through these interactions. This approach eliminates reliance on classical subroutines and achieve significant computational speedups. To the best of our knowledge, our work is the first attempt to implement a complete quantum framework for reinforcement learning by integrating the quantum realization of the classical Markov Decision Process (MDP).
@@ III.2 Quantum computing
Quantum computing (QC) combines quantum mechanics with computer science to enable advanced computation and information processing [39]. In QC, information is represented using qubits, which can exist in superposition states. This unique property allows qubits to process vast amounts of information simultaneously. The quantum computation process consists of three primary steps. The first step is state initialization, where qubits are typically prepared into uniform superposition. The second step involves the application of quantum gates to perform the computation. These gates are represented by unitary matrices, ensuring that transformations applied to quantum states are reversible and that the total probability is preserved [40]. The final step is measurement, where qubits collapse from their quantum superposition into classical states, producing probabilistic outcomes that represent the results of the computation. These steps collectively enable the construction of complex quantum algorithms capable of solving advanced problems such as Shor’s algorithm for integer factorization [41, 42, 43] and Grover’s algorithm for searching unsorted datasets [12]. In this work, Grover’s algorithm is used as a quantum subroutine to search through possible trajectories and identify those that yield the highest return. Fig. 2 presents a simple example of a Grover’s algorithm quantum circuit for 2 qubits, which searches for the target state |$\ket{11}$. This circuit illustrates the key steps of the algorithm, which include initialization, computation (oracle marking and amplitude amplification), and measurement. It conceptually demonstrates how Grover’s mechanism drives our quantum trajectory search process presented in Sec. V.
Figure 2: Quantum circuit for Grover’s algorithm on 2 qubits, searching for the target state |$\ket{11}$ generated by Qiskit [44]. The measurement was performed using IBM quantum processor (ibm_brisbane, version: 1.1.62, processor type: Eagle r3, qubits: 127) . Output distribution is displayed on the right, showing the search state |$\ket{11}$ with the highest count.
@@ IV.1 Quantum implementation of Markov Decision Process
In reinforcement learning, the agent interacts with the environment within a Markov Decision Process (MDP) framework. The agent selects actions based on a policy, which maps states to actions, while the environment determines state transitions and provides rewards. This interaction is cyclic: the agent observes a state, takes an action, receives a reward, and updates its policy to maximize cumulative rewards over time. Key MDP components include states, actions, a state transition function, and a reward function. For the MDP framework, we consider a simple stochastic environment with four states $(s_{0},s_{1},s_{2},s_{3})$ and two actions $(a_{0},a_{1})$, where the agent moves from one state to another with certain probabilities depending on the chosen action. For example, if the agent is in state $s_{0}$ and selects action $a_{1}$, there is some probability that it will move to $s_{1}$, and some probability it will remain in $s_{0}$ or will move to another state. This probability of transition from one state to another is determined by the state transition function. The reward function assigns a numerical value to each transition, representing the benefit of moving from one state to another. For instance, transitioning to state $s_{1}$ might yield a reward $r_{1}$. The goal of the agent is to find the optimal policy, a strategy that determines which action to take in each state. The optimal policy aims to maximize the total reward the agent can collect over time. Classical Q-learning solves this problem by allowing the agent to learn from its interactions with the environment. The agent explores the states, takes actions, and observes the resulting rewards and transitions. It builds a table, known as a Q-table, that estimates the expected cumulative reward for each state-action pair. By balancing exploration (trying new actions) and exploitation (choosing actions with the highest known rewards), the agent identifies the optimal actions for each state.
@@ IV.1.1 Agent and environment on quantum computer
As we are implementing MDP in the quantum domain, these states and actions can be prepared into superposition by using Hadamard transformation. This process can initialize qubits as a uniform combination of quantum samples for all states and actions within their respective spaces. For the state space S, we define the initial ground state of each state register as |$\ket{0_{s}}=\ket{0}$. Applying the Hadamard transformation to these registers generates a uniform superposition of quantum states, as formulated in Eq. (2).
$H(\ket{0_{s}}\otimes\ket{0_{s}})=\sum_{n=0}^{N-1}c_{n}\ket{s_{n}}.$ (2)
The output quantum state is a uniform superposition of the basis states, |$\ket{s_{0}},\ket{s_{1}},...,\ket{s_{n}}$, within its state space S. N represents the total number of states in that space. Each coefficient $c_{n}$ represents the amplitude of the corresponding basis state and is uniformly set to $\frac{1}{\sqrt{N}}$, ensuring that the resulting quantum state is properly normalized. This satisfies the normalization condition $\sum_{n=0}^{N-1}|c_{n}|^{2}=1$. In the case of a 2-qubit state space where $N=4$, each amplitude becomes $c_{n}=\frac{1}{2}$.
Similarly, the set of possible actions is encoded using the Hadamard operator to create a uniform superposition of all actions that belong to the action space A, as formalized in Eq. (3). The initial ground state of the action register is defined as |$\ket{0_{a}}=\ket{0}$, and applying the Hadamard transformation results in a uniform superposition where each amplitude is $c_{n}=\frac{1}{\sqrt{2}}$, satisfying the normalization condition within the action space.
$H\ket{0_{a}}=\sum_{n=0}^{N-1}c_{n}\ket{a_{n}}.$ (3)
@@
After preparing the state S and action A into uniform superposition quantum samples, we then consider the state transition function, which defines the likelihood of transitioning from one state to another given a specific action. In a classical MDP, the state transition function $P(s^{\prime}|s,a)$ determines the probability of transitioning to next state ${s^{\prime}}$ given the current state $s$ and action $a$. In a quantum circuit, these probabilities are encoded in the amplitudes of quantum states, where the squared magnitudes of the amplitudes correspond to probabilities. This is achieved by applying $R_{y}(\theta)$ to an ancillary qubit initialized in the |$\ket{0}$ state, producing a superposition in which the probabilities of measuring |$\ket{0}$ and |$\ket{1}$ are $\cos^{2}\left(\frac{\theta}{2}\right)\text{ and }\sin^{2}\left(\frac{\theta}{2}\right)$, respectively. To encode classical transition probabilities $P(s^{\prime}|s,a)$, the rotation angle $\theta$ is calculated as $\theta=2\arcsin(\sqrt{P(s^{\prime}|s,a)})$. In our work, we use multi-controlled $R_{y}(\theta)$ gates to apply these rotations conditionally based on the current state |$\ket{s}$ and action |$\ket{a}$. The rotation is applied only when the control qubits match a specific state-action pair. Equation (6) is a mathematically representation of this quantum state transition process where the controlled $R_{y}(\theta)$ gates apply a rotation to an ancillary qubit |$\ket{0_{s^{\prime}}}$ by an angle $\theta$, depending on the state-action pair, thereby determining the next state, |$\ket{s^{\prime}}\in S$.
@@ IV.2 Agent-environment quantum interactions over multiple time steps
To extend the QMDP framework over T time steps, we need to account for the quantum state transition and the calculation of cumulative rewards. At each time step t, the agent receives a quantum state |$\ket{s_{t}}$, performs an action |$\ket{a_{t}}$, and the environment responds by transitioning to a new state |$\ket{s^{\prime}_{t}}$ while providing a reward |$\ket{r_{t}}$. This process is repeated over T time steps, ensuring that the next state |$\ket{s^{\prime}_{t}}$ at t becomes the current state |$\ket{s_{t}}$ for the subsequent time step.
@@
Below is a description of how this process unfolds over $T$ time steps, including the corresponding quantum operators that govern state transitions and rewards. The approach builds on the single interaction discussed earlier and extends it to a sequence of interactions. Initially at time step ${t=0}$, the agent starts at state |$\ket{s_{0}}$, takes an action |$\ket{a_{0}}$, transition to the next state |$\ket{s^{\prime}_{0}}$ and receive a reward |$\ket{r_{0}}$. The next state |$\ket{s^{\prime}_{0}}$ propagates to the subsequent time step. At time step ${t=1}$, the agent observes the new state |$\ket{s_{1}}$ generated at ${t=0}$, selects a new action |$\ket{a_{1}}$, and the environment transitions to the next state |$\ket{s^{\prime}_{1}}$ and provides the reward |$\ket{r_{1}}$. Once again, the next state |$\ket{s^{\prime}_{1}}$ becomes the current state for the subsequent time step ${t=2}$. In this way, the interaction process will occur for $T$ time steps. The quantum state distribution at ${t=1}$ can be expressed as in Eq. (9), where $N$ represents the total number of trajectories in that quantum interaction. The terms $c_{s^{\prime}_{0},r_{0}|s_{0},a_{0}}$ and $c_{s^{\prime}_{1},r_{1}|s_{1},a_{1}}$ are the amplitudes indicating that the system transitions to the next state |$\ket{s^{\prime}}$ with reward |$\ket{r}$, given current state |$\ket{s}$ and action |$\ket{a}$ at time step ${t=0}$ and ${t=1}$, respectively.
@@
For the $T$ time steps of interaction, each quantum sample can be represented as a string of qubits encoding the possible transitions from ${t=0}$ to ${t=T-1}$, formed by concatenating state-action-next state-reward sequences. This is obtained by applying a sequence of individual interactions. The complete quantum distribution of the agent-environment interaction at time step $T$ can be summarized as Eq. (10).
|$\ket{\phi^{(T)}}=\sum_{n=1}^{N}\bigg{[}\sum_{t=0}^{T-1}c_{s^{\prime}_{t},r_{t}|s_{t},a_{t}}\prod_{t=0}^{T-1}\ket{s_{t}}\ket{a_{t}}\ket{s^{\prime}_{t}}\ket{r_{t}}\bigg{]}^{n}.$ (10)
The entire quantum state after $T$ time steps is a quantum superposition over all possible trajectories $N$, where each trajectory consists of state, action, next state and reward sequences from ${t=0}$ to ${t=T-1}$. The amplitudes $\sum_{t=0}^{T-1}c_{s^{\prime}_{t},r_{t}|s_{t},a_{t}}$ define the likelihood of each trajectory based on the agent’s actions and environment’s responses up to time step $T$.
@@ IV.2.1 Return calculation by quantum arithmetic
In our work, this classical concept is mapped into the quantum domain, where the return is encoded as a quantum state |$\ket{g}$ in a Hilbert space $G$. This space represents a quantum system with a sufficient number of qubits to encode every potential return value. Initially, the quantum state |$\ket{g}$ is prepared in its ground state |$\ket{0_{g}}$. To compute return, the quantum operation $U_{G}$ is applied to the reward qubits |$\ket{r_{t}}$ for each time step. This operation executes a quantum addition that performs sequential bit-wise addition of each reward register |$\ket{r_{t}}$ and stores the result in the return registers |$\ket{g}$. This operation propagates carry bits through the circuit, similar to a classical addition. For each reward register |$\ket{r_{t}}$, a series of controlled-NOT and Toffoli gates are applied to implement bit-wise addition into the return registers. Conceptually, this operation accumulates the discounted sum of rewards across T time steps. An abstract representation of this transformation is given in Eq. (11), where $\gamma\in[0,1]$ is a fixed discount factor, $T$ is the total number of time steps, and |$\ket{r_{t}}$ represents the reward at time step $t$.
$U_{G}:=\sum_{t=0}^{T-1}\gamma\ket{g}\bra{r_{t}}.$ | . (11)
@@
When applying the return operation $U_{G}$ to the Quantum Markov Decision Process (QMDP) over $T$ time steps, the goal is to compute the return for each possible trajectory by summing the discounted rewards across those time steps. The final quantum state over these interactions can be expressed as Eq. (12).
@@
where N is the total number of possible quantum trajectories for the agent-environment interactions over $T$ time steps, |$\ket{g}$ represents the return for each trajectory, which sums the discounted rewards over the $T$ time steps. Therefore, the complete quantum state |$\ket{\phi^{(T)}}$ contains not only the state, action, next state and reward sequences but also the return for each trajectory. This enables the QMDP to evaluate the performance of all possible sequences of interactions efficiently by computing the return across multiple trajectories simultaneously.
The quantum state |$\ket{\phi^{(T)}}$, discussed in Eq. (12), is not human-interpretable classical information. The coefficients $c_{s^{\prime}_{t},r_{t}|s_{t},a_{t}}$ that define the amplitude of the quantum state cannot be accessed directly due to the nature of quantum systems. The only way to obtain the information from the quantum state |$\ket{\phi^{(T)}}$ is by measurement. Measurement is a probabilistic process that results a specific quantum state with a likelihood determined by $|c_{s^{\prime}_{t},r_{t}|s_{t},a_{t}}|^{2}$ Therefore, the quantum state |$\ket{\phi^{(T)}}$ can be regarded as quantum encoding of probability distributions over the possible trajectories of the agent-environment interactions across $T$ time steps.